OpenStudy (anonymous):

In the triangle below, b = _____. If necessary, round your answer to two decimal places. Angle A = 33.7° Angle B = ?? Angle C = 26.4° Side a = 24 Side b = ?? Side c = ?? All I know is how to find Angle B I don't know what to do after that. 33.7° + 26.4° = 60.1° 180° - 60.1° = 119.9° So Angle B = 119.9°

OpenStudy (anonymous):

law of sines

OpenStudy (anonymous):

What is that?

OpenStudy (anonymous):

do you know trigonometry?

OpenStudy (anonymous):

Some. Is it\[\frac{ a }{ \sin A }=\frac{ b }{ \sin B }=\frac{ c }{ \sin C }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So \[\frac{ 24 }{ 33.7° }=\frac{ b }{ 119.9° }=\frac{ c }{ 26.4 }\] Now what do I do?

OpenStudy (anonymous):

?